OpenStudy (anonymous):

f(x)=sinx, x0=Pi/3 f'(x)= y=

OpenStudy (anonymous):

you looking for the derivative?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

does it help to know that the derivative of sine is cosine?

OpenStudy (anonymous):

or are you supposed to be doing this by hand without knowing that?

OpenStudy (anonymous):

i know that but is te solution?

OpenStudy (anonymous):

what i need to do with Pi/3 ?

OpenStudy (anonymous):

\[f(x)=\sin(x)\\f'(x)=\cos(x)\\ \cos(\frac{\pi}{3})=\frac{1}{2}\]

OpenStudy (anonymous):

not really sure what the question is asking though

OpenStudy (anonymous):

thanks, may be is correct